OpenStudy (anonymous):

Use each of the numbers 2,3,4,and 6 once to make each equation correct. 11. ( - ) + = 11. 12. times - ( / ) = 6. 13. + ( times ) times = 30. 14. / + times = 20.

OpenStudy (agent47):

how can u use all 2,3,4 and 6 in equation 11 if you only have 3 spots

OpenStudy (swissgirl):

Nooooo like in every specific equation u can only use a number once. I am assuming

OpenStudy (anonymous):

I mean like those blank spaces are for those answers.......... example: 2,3,4, and 6

OpenStudy (swissgirl):

idk ( - )+ =11 is impossible with ur options of 2,3,4 and 6

OpenStudy (anonymous):

I might be wrong through

OpenStudy (swissgirl):

(3-2)+4+6=11

OpenStudy (anonymous):

Thanks Swissgirl :D

OpenStudy (swissgirl):

well t was agents answer but secondly we only solved the first one

OpenStudy (anonymous):

oh okay, and do u know how to solve the others?

OpenStudy (swissgirl):

ok second one is : 4 times 2 - (6/3) = 6

OpenStudy (swissgirl):

third one is 6+(2*3)*4=30

OpenStudy (swissgirl):

Fourth one is 4/2 + 6*3 = 20

OpenStudy (anonymous):

wait, the first one is (3 - 2) + 4 times 6 = 6

OpenStudy (anonymous):

I think

OpenStudy (swissgirl):

that is incorrect

OpenStudy (swissgirl):

the answers we gave u r the correct ones

OpenStudy (anonymous):

U mean the one u wrote or mine?

OpenStudy (swissgirl):

The one I wrote.

OpenStudy (anonymous):

oh

OpenStudy (swissgirl):

I gave u first, second, third and fourth, they go in order as u have asked them

OpenStudy (anonymous):

okay then, thanx :)